Determine the present value of a 20-year GIC with an interest rate of 7.9% a compounded quarterly, if the future value is $11100
A) 150.21 B) 2426.02 C) 2321.81 D) 1387.55

Answer by Theo(13342)

i get selection C.

the formula to use is f = p * (1 + r) ^ n

f is the future value
p is the present value
r is the interest rate per time period #####
##### time period is in quarters therefore take the annual interest rate and divide it by 4 to get the quarterly interest rate (not the %).
n is the number of time periods #####
##### time period is in quarters therefore take the number of years and multiply them by 4 to get the number of quarters.

in your problem:

f = 11100
p = what you want to find
r = 7.9% / 100 = .079 / 4 = .01975
n = 20 * 4 = 80

formula becomes 11100 = p * (1.01975) ^ 80
solve for p to get p = 11100 / ((1.01975) ^ 80)
your result is p = 2321.805628

that's selection C.